arly17:

If a running system has a total change in the heat of 315 joules, and it is running at a temperature of 350 K, what is the entropy? A. 1.54 joules/kelvin B. 0.847 joules/kelvin C. 1.22 joules/kelvin D. 0.900 joules/kelvin E. 0.762 joules/kelvin

Florisalreadytaken:

\[ \text{entropy change} = \frac{ \text{change in heat}}{temperature} \Rightarrow \frac{ \text{315 joules}}{350 kelvin} \] \[ \text{entropy change} = 0.90 \ \frac{joules}{kelvin} \] which option would that be?*